# Revolutionizing Mixed Precision Quantization: Towards Training-free Automatic Proxy Discovery via Large Language Models

### v1

@@ Abstract
Mixed-Precision Quantization (MPQ) liberates the Deep Neural Networks (DNNs) from the Out-Of-Memory (OOM) bottleneck, which garnered increasing research attention. However, conventional methods either searched from costly differentiable optimization, which is neither efficient nor flexible, or learned a quantized DNN from the proxy (i.e., HAWQ) manually designed by human experts, which is labor-intensive and requires huge expert knowledge. Can we design a proxy without involving any human experts and training? In this paper, we provide an affirmative answer by proposing a novel Large Language Models (LLMs)-driven Training-free Automatic Proxy (dubbed, TAP) discovery framework, which reforms the design paradigm of MPQ by utilizing LLMs to find superior TAP tailored for MPQ, automatically. In addition, to bridge the gap between black-box LLMs and the tough MPQ task, we ingeniously propose simple Direct Policy Optimization (DPO) based reinforcement learning to enhance LLMs’s reasoning by optimizing prompts, which can construct positive feedback loop between the LLM and the MPQ task, enabling LLMs to generate better TAP in the next evolution. Extensive experiments on mainstream benchmarks demonstrate that TAP achieves state-of-the-art performance. Finally, we truly believe that our TAP will significantly contribute to the MPQ community by providing a new perspective on LLM-driven design algorithms.
@@ 1 Introduction
Deep neural networks have become indispensable in modern vision applications. However, when deployed on extremely resource-limited devices such as MCUs and tiny NPUs, they often encounter severe Out-Of-Memory (OOM) issues due to the gap between model complexity and the tiny on-chip memory. This tension motivates the development of efficient compression techniques. Although fixed-precision quantization (FPQ) [choi2018pact] provides a straightforward solution, its uniform bit assignment usually leads to unsatisfactory accuracy under strict hardware budgets. Mixed-precision quantization (MPQ) [wang2019haq, dong2019hawq, yao2021hawq], in contrast, adjusts bit-widths according to layer sensitivity and thus achieves a better balance between accuracy and efficiency. Existing differentiable MPQ methods [cai2020rethinking, zhang2021differentiable, chu2021mixed, yang2021fracbits] further attempt to learn bit allocations automatically, yet their large computational cost severely limits practical deployment (Table 1). As a response, training-free MPQ methods (e.g., HAWQ [dong2019hawq], HAWQ-V2 [6], OMPQ [ma2023ompq]) avoid huge training costs, although they still require considerable manual effort.
Figure 1: TAP vs. Peer competitors in terms of accuracy, generalization, and costs.
@@
EdMIPS-Img [cai2020rethinking] 65.9 $\ast$$\ast$ 34.7 9.5
EdMIPS-C [cai2020rethinking] 59.1 $\ast$$\ast$ 7.4 0.6
@@
HAWQ [dong2019hawq] $\begin{aligned} \max_{i}\{\lambda_{i}(H)\}\end{aligned}$ - 2500
@@
Contributions. In this work, we aim to analyze and address the above limitations. We first revisit the design principles of existing proxies and empirically verify their drawbacks in terms of expert dependency, calibration cost and convergence behavior (Section 2). Motivated by the advancement of LLMs in generating structured knowledge and reasoning [radford2018improving, guo2025deepseek], we propose TAP, the first framework that automatically constructs training-free MPQ proxies. At the beginning, we examine a simple prompting strategy. However, similar to the observations in other automated tasks, directly prompting the LLM yields unstable layer-wise bit allocations. A closer analysis reveals that this is largely due to the absence of any feedback signal that connects the generated proxy to its MPQ performance. To overcome this issue, we draw inspiration from the reasoning behaviors observed in Qwen3 [qwen3] and DeepSeek [guo2025deepseek], and introduce a lightweight Direct Policy Optimization (DPO) [rafailov2024directpreferenceoptimizationlanguage] reinforcement learning procedure. This introduces task-aware feedback into the prompt evolution, encouraging the LLM to produce intermediate reasoning steps before generating the final proxy. Thereby, the proxy quality improves steadily through iterative refinement. Our contributions are summarized as follows:
  • A new proxy-design paradigm for MPQ. We present TAP, an LLM-driven framework that automatically discovers training-free MPQ proxies without relying on handcrafted rules, offering a fresh perspective on MPQ design.
  • Enhanced LLM reasoning through DPO. We identify the main weakness of naive prompting and introduce a DPO-based strategy that provides task-aware feedback, leading to more reliable Chain-of-Thought style reasoning and stronger accuracy–efficiency performance.
@@ 2.2 Observations and Motivations
As shown in Table 2, we can clearly observe that existing proxies (i.e., HAWQ, OMPQ) for training-free MPQ still depend on hand-crafted rules (e.g., Hessian traces or activation–weight statistics), which results in huge labor costs. Moreover, those methods suffer from large calibration sets, such as HAWQ-V2 requiring 8,1928{,}192 samples and $50$ iterations, and HAWQ over 2,5002{,}500 updates. Even OMPQ relies on manual patterns, making these methods costly and sensitive to expert design. For a new task or dataset, MPQ proxies face a dilemma: they must either be re-searched from scratch, which is computationally expensive and inflexible, or transferred from proxies designed for other tasks, which risks sub-optimal quantization performance due to task mismatch. Motivated by the reasoning abilities of LLMs, we investigate whether they can automatically generate effective training-free proxies for MPQ. Using a simple prompt (App. A), LLMs can produce proxy candidates, but naive prompting yields inconsistent performance due to the lack of task-aware feedback. To overcome this, we propose a Direct Policy Optimization (DPO) based reinforcement learning framework, inspired by Chain-of-Thought (CoT) prompting in LLMs (i.e., Deepseek). This establishes a feedback loop between the LLM and the MPQ task, allowing iterative refinement of proxies with lightweight reward signals. As a result, LLM-driven proxy generation reduces dependence on experts and calibration, while producing more accurate and hardware-aware proxies, offering a scalable and efficient paradigm for training-free MPQ.
@@ 3.1 Overview of TAP
Figure 2: Overview of TAP. Input the initial prompt first. The LLM generates an initial proxy population, where each proxy must include inference logic, executable code, and bit-width allocation. Candidate proxies are validated on benchmarks like ImageNet-1k, with quantization accuracy and computational cost calculated to produce an adaptability score. Using this score, construct proxy preference data pairs, optimize LLM parameters via the DPO module to guide subsequent proxy generation, merge the current and candidate populations, and filter by adaptability score to maintain a fixed size.
To achieve the goal of designing zero-cost proxies for mixed-precision quantization tasks, the TAP framework leverages large language models to automatically generate proxies by optimizing natural language descriptions and corresponding code. Furthermore, we propose an evolution scheduler based on Direct Preference Optimization (DPO), which optimizes the direction of prompt generation by learning strategies aligned with human preferences (or task performance preferences). This aims to maximize the correlation between proxies and the final model performance while avoiding issues such as complex reward function design and unstable training in traditional reinforcement learning. The overall architecture of the TAP framework is illustrated in Fig. 2, consisting of three core components:
@@
RL Evolution Scheduler. To enable the proxy evolutionary strategy to be learnable and efficiently converge to optimal zero-cost proxies, the TAP framework incorporates a lightweight Direct Preference Optimization (DPO) module that serves as an evolutionary decision-maker. This DPO module does not require explicit reward function design; instead, it takes the "preference data pairs" (constructed by comparing the fitness scores of different proxies returned by the evaluator) as supervision signals. It learns a prompt generation strategy that makes the generated candidate proxies more aligned with the task’s preference for "high performance and efficiency," thereby co-optimizing both the evolving set of zero-cost proxies and the evolutionary strategy itself.
@@ 3.2 Automatic Proxy Discovery for Mixed-Precision Quantization
Proxy Candidate Generator. Let $\mathcal{P}$ denote the set of valid proxies that satisfy the fixed input-output contract for mixed-precision quantization tasks. Specifically, each proxy must generate sensitivity scores for convolutional layer channels and linear layers of neural networks—these scores directly guide the bit-width allocation process. At time step $t$, given the current proxy population $\mathcal{P}_{t}\ (\mathcal{P}_{t}\subseteq\mathcal{P})$, the Large Language Model (LLM) receives two types of inputs simultaneously: 1. A structured prompt $\pi\in\Pi$ specifying the operation type (initialization, mutation, crossover); 2. A bounded context window containing existing proxies, including their natural language reasoning logic, executable code for sensitivity scoring, and historical bit-allocation performance.
This input pair induces a context-conditioned distribution:
$\mu_{\pi,\mathcal{C}_{t}}=P_{\mathcal{L}}\big(f\mid\pi,\mathcal{C}_{t}\big)$ (4)
@@
$\mathcal{F}=\bigcup_{t\in\mathbb{N}}\bigcup_{\pi\in\Pi}\bigcup_{\mathcal{C}\subseteq\mathcal{P}_{t}}\mathrm{supp}\big(\mu_{\pi,\mathcal{C}}\big)$ (5)
@@
Here, $\text{supp}(\mu_{\pi,\mathcal{C}})$ denotes the support set of the distribution $\mu_{\pi,\mathcal{C}}$, and $\mathcal{F}\subseteq\mathcal{P}$. This structure allows the generator to reuse effective sensitivity scoring patterns and bit-allocation heuristics discovered in previous iterations.
@@
For mixed-precision quantization tasks, three categories of context-conditioned input pairs are designed, corresponding to different proxy generation logics: - Initialization $\mu_{\text{init},\emptyset}$: Provides task descriptions (sensitivity scoring for convolutional/linear layers, bit-width allocation), input-output contracts (input: network weights/activations; output: channel-level/layer-level sensitivity scores), and prior knowledge (e.g., "sensitivity scores are positively correlated with quantization error—higher scores require more bits"). - Mutation $\mu_{\text{mut},f}$: Optimizes the sensitivity scoring logic (e.g., adjusting statistical metrics for convolutional channel importance) or bit-allocation function parameters based on a single proxy $f$, enabling fine-grained optimization. - Crossover $\mu_{\text{cross},f_{1},f_{2}}$: Fuses complementary components of two parent proxies (e.g., inheriting convolutional layer scoring logic from $f_{1}$ and linear layer scoring logic from $f_{2}$) to enhance the robustness of proxies across different network layers.
@@
Fitness Evaluator. For a benchmark set $\mathcal{A}$ composed of neural networks (e.g., ResNet, ViT) and their ground-truth performance on the ImageNet-1k dataset, the fitness of a proxy $f$ is quantified by two core metrics: 1. The quality of sensitivity scores (measured by the Spearman correlation between predicted sensitivity and actual quantization error); 2. The effectiveness of bit-width allocation (measured by the top-1 accuracy of the quantized model on ImageNet-1k, aligned with the target compression rate).
The fitness function is defined as:
$\phi(f)=\alpha\cdot\rho_{\text{sens}}+(1-\alpha)\cdot\mathrm{Acc}_{\text{quant}}$ (6)
@@
where: $\rho_{\text{sens}}$ is the Spearman correlation between the sensitivity scores generated by the proxy and the empirical quantization error (smaller error indicates higher sensitivity); $\text{Acc}_{\text{quant}}$ is the top-1 accuracy of the quantized model on ImageNet-1k using the bit-width configuration derived from the proxy; $\alpha\in[0,1]$ balances the quality of sensitivity scoring and quantization accuracy, and $\beta\geq 0$ controls computational efficiency.
@@
DPO aims to optimize a prompt generation strategy aligned with task preferences (high fitness), enabling LLMs to produce candidate proxies that better satisfy “strong sensitivity-accuracy correlation and low computational cost". Its core formula, based on the probability ratio between the optimized model and reference model, is:
$\log\frac{P(\pi)}{P_{0}(\pi)}\geq\lambda\cdot\mathbb{E}\Big[\log\frac{P(f_{+}\mid\pi)}{P(f_{-}\mid\pi)}\Big]$ (7)
where $P(\cdot)$ denotes the optimized LLM’s probability distribution, and $P_{0}(\cdot)$ the initial reference LLM (without preference adjustment); - $\pi$ is the proxy-generation prompt (context window $\mathcal{C}_{t}$ and mixed-precision quantization task description $\mathcal{T}$ are implicitly included in model inputs); - $\lambda$ is a regularization coefficient balancing the magnitude and stability of strategy updates; - $\mathbb{E}$ represents the expectation over preference data pairs $(f_{+},f_{-})$ from set $\mathcal{D}$ (sorted by fitness scores), where $f_{+}$ and $f_{-}$ denote preferred and dispreferred proxies, respectively.
By minimizing the above objective function, DPO can iteratively optimize the prompt response logic of the LLM: After updating $\mathcal{L}_{\theta}$ based on the current preference data pairs at generation $t$, the LLM at the next generation ($t+1$) will generate proxies with better fitness. This process simultaneously optimizes both the proxy population itself and the evolution strategy for proxy generation, forming a positive feedback loop of “prompt strategy-proxy quality".
Algorithm 1 Evolutionary Framework of TAP (DPO-Optimized Version)
1:Reference LLM $\mathcal{L}_{\text{ref}}$, Optimizable LLM $\mathcal{L}_{\theta}$, ImageNet-1k benchmark $\mathcal{B}$, Network set $\mathcal{A}$, Population size NN, Max generations $T_{\text{max}}$, Target compression rate $\zeta$, DPO regularization coefficient $\lambda$
@@
5:  Sample context window $\mathcal{C}_{t}\subseteq\mathcal{K}$, sample operation type $\text{op}\in\{\text{mutation},\text{crossover}\}$
6:  $\mathcal{P}_{t}^{\prime}\leftarrow\mathcal{L}_{\theta}(\mu_{\text{op},\mathcal{C}_{t}})$ (generate candidate proxies for mixed-precision quantization)
7:  for each $f\in\mathcal{P}_{t}^{\prime}$ do
8:   Compute sensitivity scores of convolutional channels/linear layers for networks in $\mathcal{A}$
9:   Allocate bits to networks via $f$’s bit-allocation function and $\zeta$
10:   Quantize networks and evaluate accuracy on ImageNet-1k
11:   Compute $\phi(f)$ using the fitness function in Section 3.2
12:  end for
13:  Construct preference dataset $\mathcal{D}$: Sort $\mathcal{P}_{t}\cup\mathcal{P}_{t}^{\prime}$ by $\phi(f)$, generate $(f_{\text{prefer}},f_{\text{disprefer}})$ pairs
14:  DPO strategy update: Minimize the core optimization formula in Section 3.2, update parameters $\theta$ of $\mathcal{L}_{\theta}$
15:  Update context library $\mathcal{K}$: Retain top-10% high-fitness proxies in $\mathcal{P}_{t}\cup\mathcal{P}_{t}^{\prime}$
16:  $\mathcal{P}_{t+1}\leftarrow$ Top-NN proxies from $\mathcal{P}_{t}\cup\mathcal{P}_{t}^{\prime}$ (sorted by $\phi(f)$)
17:end forreturn $\arg\max_{f\in\bigcup_{t=0}^{T_{\text{max}}}\mathcal{P}_{t}}\phi(f)$ (optimal TAP proxy for mixed-precision quantization)

### v2

@@ Abstract
Mixed-Precision Quantization (MPQ) liberates the Deep Neural Networks (DNNs) from the Out-Of-Memory (OOM) bottleneck, which garnered increasing research attention. However, conventional methods either searched from costly differentiable optimization, which is neither efficient nor flexible, or learned a quantized DNN from the proxy (i.e., HAWQ) manually designed by human experts, which is labor-intensive and requires huge expert knowledge. Can we design a proxy without involving any human experts and training? In this paper, we provide an affirmative answer by proposing a novel Large Language Models (LLMs)-driven Training-free Automatic Proxy (dubbed, TAP) discovery framework, which reforms the design paradigm of MPQ by utilizing LLMs and evolutionary search strategies to find superior TAP tailored for MPQ, automatically. In addition, to bridge the gap between black-box LLMs and the tough MPQ task, we introduce a lightweight Direct Preference Optimization (DPO)–based strategy controller that dynamically reweights the selection probabilities of the three prompt templates for evolutionary search strategies according to fitness signals, without fine-tuning the LLM. This builds a task-aware feedback loop that improves proxy generation across evolutions. Extensive experiments on mainstream benchmarks demonstrate that TAP achieves state-of-the-art performance. Finally, we truly believe that our TAP will significantly contribute to the MPQ community by providing a new perspective on LLM-driven design algorithms.
@@ 1 Introduction
Deep neural networks have become indispensable in modern vision applications. However, when deployed on extremely resource-limited devices such as MCUs and tiny NPUs, they often encounter severe Out-Of-Memory (OOM) issues due to the gap between model complexity and the tiny on-chip memory. This tension motivates the development of efficient compression techniques. Although fixed-precision quantization (FPQ) [1] provides a straightforward solution, its uniform bit assignment usually leads to unsatisfactory accuracy under strict hardware budgets. Mixed-precision quantization (MPQ) [2, 3, 4], in contrast, adjusts bit-widths according to layer sensitivity and thus achieves a better balance between accuracy and efficiency. Existing differentiable MPQ methods [5, 6, 7, 8] further attempt to learn bit allocations automatically, yet their large computational cost severely limits practical deployment (Table 1). As a response, training-free MPQ methods (e.g., HAWQ [3], HAWQ-V2 [9], OMPQ [10]) avoid huge training costs, although they still require considerable manual effort.
@@
EdMIPS-Img [5] 65.9 $\ast$$\ast$ 34.7 9.5
EdMIPS-C [5] 59.1 $\ast$$\ast$ 7.4 0.6
@@
HAWQ [3] $\begin{aligned} \max_{i}\{\lambda_{i}(H)\}\end{aligned}$ - 2500
@@
Contributions. In this work, we aim to analyze and address the above limitations. We first revisit the design principles of existing proxies and empirically verify their drawbacks in terms of expert dependency, calibration cost and convergence behavior (Section 2). Motivated by the advancement of LLMs in generating structured knowledge and reasoning [11, 12], we propose TAP, the first framework that automatically constructs training-free MPQ proxies. At the beginning, we examine a simple prompting strategy. However, similar to the observations in other automated tasks, directly prompting the LLM yields unstable layer-wise bit allocations. A closer analysis reveals that this is largely due to the absence of any feedback signal that connects the generated proxy to its MPQ performance. To overcome this issue, we draw inspiration from the reasoning behaviors observed in Qwen3 [13] and DeepSeek [12], and introduce a lightweight Direct Policy Optimization (DPO) [14] reinforcement learning procedure. This introduces task-aware feedback into the prompt evolution, encouraging the LLM to produce intermediate reasoning steps before generating the final proxy. Thereby, the proxy quality improves steadily through iterative refinement. Our contributions are summarized as follows:
  • A new proxy-design paradigm for MPQ. We present TAP, an LLM-driven framework that adopts evolutionary search strategies and automatically discovers training-free proxies for mixed-precision quantization without relying on handcrafted rules, thus offering a fresh perspective on the design of mixed-precision quantization.
  • DPO as a template selector. We identify the core weakness of naive prompting strategies and introduce a direct preference optimization (DPO)-based strategy controller that reallocates selection probabilities for three prompt templates according to fitness signals—without updating large language model (LLM) parameters—resulting in more reliable reasoning during proxy generation and superior integrated accuracy–efficiency performance.
@@ 2.2 Observations and Motivations
As shown in Table 2, we can clearly observe that existing proxies (i.e., HAWQ, OMPQ) for training-free MPQ still depend on hand-crafted rules (e.g., Hessian traces or activation–weight statistics), which results in huge labor costs. Moreover, those methods suffer from large calibration sets, such as HAWQ-V2 requiring 8,1928{,}192 samples and $50$ iterations, and HAWQ over 2,5002{,}500 updates. Even OMPQ relies on manual patterns, making these methods costly and sensitive to expert design. For a new task or dataset, MPQ proxies face a dilemma: they must either be re-searched from scratch, which is computationally expensive and inflexible, or transferred from proxies designed for other tasks, which risks sub-optimal quantization performance due to task mismatch. Motivated by the reasoning abilities of LLMs, we investigate whether they can automatically generate effective training-free proxies for MPQ. Using a simple prompt (App. A), LLMs can produce proxy candidates, but naive prompting yields inconsistent performance due to the lack of task-aware feedback. To overcome this, we adopt a Direct Preference Optimization (DPO)–based strategy controller that reweights three prompt templates using fitness feedback, without fine-tuning the LLM. This establishes a task-aware feedback loop that iteratively refines proxy quality with lightweight signals. As a result, LLM-driven proxy generation reduces dependence on experts and calibration, while producing more accurate and hardware-aware proxies, offering a scalable and efficient paradigm for training-free MPQ. Importantly, the TAP proxy metric presented in Table 2 is only one example among multiple optimal proxy-metric formulas generated by the large language model. This particular formula reorganizes and quantifies existing heuristic rules using mathematical constructs such as norms, entropy, and exponential decay. By drawing on vast domain knowledge, the LLM can extract and recombine effective heuristics; the resulting algorithms inherit the rationality of human expertise while overcoming its inherent limitations.
@@ 3.1 Overview of TAP
Figure 1: Overview of TAP. Input the initial prompt first. The LLM generates an initial proxy population, where each proxy must include inference logic, executable code, and bit-width allocation. Candidate proxies are validated on benchmarks like ImageNet-1k, with quantization accuracy and computational cost calculated to produce an adaptability score. Using this score, construct proxy preference data pairs and use a DPO-based strategy controller to dynamically adjust the selection probabilities of three prompt templates (without fine-tuning the LLM) for subsequent proxy generation. Merge the current and candidate populations, and filter by adaptability score to maintain a fixed size.
To achieve the goal of designing zero-cost proxies for mixed-precision quantization tasks, the TAP framework leverages large language models and evolutionary search strategies to automatically generate proxies by optimizing natural language descriptions and their corresponding code. Furthermore, we propose an evolution scheduler based on Direct Preference Optimization (DPO), which acts as a non-parametric strategy controller that reallocates the selection probabilities of the three prompt templates for evolutionary search strategies according to fitness signals, without any fine-tuning of the large language models (LLMs). The overall architecture of the TAP framework is illustrated in Fig. 1, consisting of three core components:
@@
RL Evolution Scheduler. To enable the proxy evolutionary strategy to be learnable and efficiently converge to optimal zero-cost proxies, the TAP framework incorporates a lightweight Direct Preference Optimization (DPO) module that serves as an evolutionary decision-maker. This DPO module does not require explicit reward function design; instead, it takes the "preference data pairs" (constructed by comparing the fitness scores of different proxies returned by the evaluator) as supervision signals and updates the selection weights of three prompt templates. Importantly, the LLM parameters remain fixed throughout; DPO only adjusts template selection probabilities to make the generated candidate proxies more aligned with the task’s preference for "high performance and efficiency."
@@ 3.2 Automatic Proxy Discovery for Mixed-Precision Quantization
Proxy Candidate Generator. Let $\mathcal{P}$ denote the set of valid proxies that satisfy the fixed input-output contract for mixed-precision quantization tasks. Specifically, each proxy must generate sensitivity scores for convolutional layer channels and linear layers of neural networks—these scores directly guide the bit-width allocation process. At time step $t$, given the current proxy population $\mathcal{P}_{t}\ (\mathcal{P}_{t}\subseteq\mathcal{P})$, the Large Language Model (LLM) receives two types of inputs simultaneously: 1. A structured prompt $\pi\in\Pi$ specifying the operation type (initialization, mutation, crossover); 2. A bounded context window containing existing proxies, including their natural language reasoning logic, executable code for sensitivity scoring, and historical bit-allocation performance. This input pair induces a context-conditioned distribution:
$\mu_{\pi,\mathcal{C}_{t}}=P_{\mathcal{L}}\big(f\mid\pi,\mathcal{C}_{t}\big),$ (4)
@@
$\mathcal{F}=\bigcup_{t\in\mathbb{N}}\bigcup_{\pi\in\Pi}\bigcup_{\mathcal{C}\subseteq\mathcal{P}_{t}}\mathrm{supp}\big(\mu_{\pi,\mathcal{C}}\big),$ (5)
@@
where $\text{supp}(\mu_{\pi,\mathcal{C}})$ denotes the support set of the distribution $\mu_{\pi,\mathcal{C}}$, and $\mathcal{F}\subseteq\mathcal{P}$. This structure allows the generator to reuse effective sensitivity scoring patterns and bit-allocation heuristics discovered in previous iterations.
Search Space and Constraints. The proposed method performs an evolutionary search entirely within two coupled discrete spaces that are instantiated and navigated via prompt templates, without any first- or second-order model information: - Proxy space $\mathcal{F}$. Each candidate proxy $f\in\mathcal{F}$ is represented as a tuple $(T,C)$, where $T$ is a natural-language reasoning trace and $C$ is deterministic, executable code that produces a scalar score used to rank mixed-precision configurations. Proxies adhere to a fixed input–output contract and may only use architecture- and task-level metadata that do not require accessing internal training signals. In particular, the proxy design explicitly excludes the use of Hessian matrices, gradients, or raw weight/activation tensors. - Configuration space $\mathcal{A}$. Each configuration $a\in\mathcal{A}$ is a vector of per-layer bit-width assignments subject to hardware or compression constraints. In experiments, $\mathcal{A}$ is instantiated with standard discrete bit choices for weights and activations (e.g., weights in $\{2,4,8\}$ and activations fixed at $8$ bits or chosen from $\{2,4,8\}$, depending on the setting), but the framework is agnostic to the exact candidate sets.
Evolution proceeds by sampling from context-conditioned distributions $\mu_{\pi,\mathcal{C}_{t}}$ using three prompt templates (initialization, mutation, crossover), while a lightweight DPO controller reweights template-selection probabilities based on fitness signals. Crucially, the LLM parameters remain frozen and the entire search avoids any reliance on Hessian traces, gradient-based signals, or weight-tensor statistics; guidance comes solely from black-box task feedback (e.g., rank-correlation with observed performance and simple population measures), keeping the pipeline strictly training-free.
@@
For mixed-precision quantization tasks, three categories of context-conditioned input pairs are designed, corresponding to different proxy generation logics: - Initialization $\mu_{\text{init},\emptyset}$: Provides task descriptions (sensitivity scoring for convolutional/linear layers, bit-width allocation), input–output contracts (input: architecture metadata and compression targets; output: channel-level/layer-level sensitivity scores), and prior knowledge (e.g., "sensitivity scores are positively correlated with quantization error—higher scores require more bits"). - Mutation $\mu_{\text{mut},f}$: Optimizes the sensitivity scoring logic (e.g., adjusting statistical metrics for convolutional channel importance) or bit-allocation function parameters based on a single proxy $f$, enabling fine-grained optimization. - Crossover $\mu_{\text{cross},f_{1},f_{2}}$: Fuses complementary components of two parent proxies (e.g., inheriting convolutional layer scoring logic from $f_{1}$ and linear layer scoring logic from $f_{2}$) to enhance the robustness of proxies across different network layers.
@@
Fitness Evaluator. For a benchmark set $\mathcal{A}$ composed of neural networks (e.g., ResNet, ViT) and their ground-truth performance on the ImageNet-1k dataset, the fitness of a proxy $f$ is quantified by two core metrics: 1. The quality of sensitivity scores (measured by the Spearman correlation between predicted sensitivity and actual quantization error); 2. The effectiveness of bit-width allocation (measured by the top-1 accuracy of the quantized model on ImageNet-1k, aligned with the target compression rate). The fitness function is defined as:
$\phi(f)=\alpha\cdot\rho_{\text{sens}}+(1-\alpha)\cdot\mathrm{Acc}_{\text{quant}},$ (6)
@@
where $\rho_{\text{sens}}$ is the Spearman correlation between the sensitivity scores generated by the proxy and the empirical quantization error (smaller error indicates higher sensitivity); $\text{Acc}_{\text{quant}}$ is the top-1 accuracy of the quantized model on ImageNet-1k using the bit-width configuration derived from the proxy; $\alpha\in[0,1]$ balances the quality of sensitivity scoring and quantization accuracy, and $\beta\geq 0$ controls computational efficiency.
@@
DPO serves as a non-parametric controller over a set of three prompt templates $\mathcal{T}=\{T_{1},T_{2},T_{3}\}$. At generation $k$, each template $T_{t}$ produces a subset of candidate proxies whose fitness is aggregated into $\mathrm{Fitness}(T_{t})$ (e.g., average over its candidates). The controller maintains nonnegative weights $\{w_{t}^{(k)}\}_{t=1}^{3}$ and updates them by
$w_{t}^{(k+1)}\;=\;w_{t}^{(k)}\cdot\alpha\cdot\frac{\mathrm{Fitness}(T_{t})}{\mu^{(k)}},$ (7)
where $\mu^{(k)}$ is the mean fitness across all templates at generation $k$, and $\alpha>0$ controls the update scale. The selection probability of each template is then obtained by normalizing the weights:
$p_{t}^{(k+1)}\;=\;\frac{w_{t}^{(k+1)}}{\sum_{j=1}^{3}w_{j}^{(k+1)}}\,.$ (8)
Templates with higher fitness receive larger selection probabilities in subsequent generations. Crucially, the LLM parameters are never updated; DPO only adjusts template selection probabilities.
Algorithm 1 Evolutionary Framework of TAP (DPO-based Template Selection)
1:LLM $\mathcal{L}$, prompt templates $\mathcal{T}=\{T_{1},T_{2},T_{3}\}$, initial weights $w^{(0)}$ (uniform), ImageNet-1k benchmark $\mathcal{B}$, Network set $\mathcal{A}$, Population size NN, Max generations $T_{\text{max}}$, Target compression rate $\zeta$
@@
5:  Compute $p^{(t)}$ from $w^{(t)}$ using Eq. (8); sample $\text{op}\in\{\text{initialization},\text{mutation},\text{crossover}\}$
6:  Sample $T\sim\text{Categorical}(p^{(t)})$
7:  if $\text{op}=\text{initialization}$ then
8:   $\mathcal{C}_{t}\leftarrow\emptyset$; $\mathcal{P}_{t}^{\prime}\leftarrow\mathcal{L}(\mu_{\text{init},\emptyset},T)$
9:  else
10:   Sample context $\mathcal{C}_{t}\subseteq\mathcal{K}$; $\mathcal{P}_{t}^{\prime}\leftarrow\mathcal{L}(\mu_{\text{op},\mathcal{C}_{t}},T)$
11:  end if
12:  for each $f\in\mathcal{P}_{t}^{\prime}$ do
13:   Compute sensitivity scores of convolutional channels/linear layers for networks in $\mathcal{A}$
14:   Allocate bits to networks via $f$’s bit-allocation function and $\zeta$
15:   Quantize networks and evaluate accuracy on ImageNet-1k
16:   Compute $\phi(f)$ using the fitness function in Section 3.2
17:  end for
18:  Aggregate per-template fitness $\mathrm{Fitness}(T_{i})$; update $w^{(t+1)}$ via Eq. (7)
19:  Update context library $\mathcal{K}$: Retain top-10% high-fitness proxies in $\mathcal{P}_{t}\cup\mathcal{P}_{t}^{\prime}$
20:  $\mathcal{P}_{t+1}\leftarrow$ Top-NN proxies from $\mathcal{P}_{t}\cup\mathcal{P}_{t}^{\prime}$ (sorted by $\phi(f)$)
21:end forreturn $\arg\max_{f\in\bigcup_{t=0}^{T_{\text{max}}}\mathcal{P}_{t}}\phi(f)$ (optimal TAP proxy for mixed-precision quantization)